# Multi-Agent Quantum Reinforcement Learning using Evolutionary Optimization

### v3

@@ Abstract
Multi-Agent Reinforcement Learning is becoming increasingly more important in times of autonomous driving and other smart industrial applications. Simultaneously a promising new approach to Reinforcement Learning arises using the inherent properties of quantum mechanics, reducing the trainable parameters of a model significantly. However, gradient-based Multi-Agent Quantum Reinforcement Learning methods often have to struggle with barren plateaus, holding them back from matching the performance of classical approaches. We build upon an existing approach for gradient free Quantum Reinforcement Learning and propose three genetic variations with Variational Quantum Circuits for Multi-Agent Reinforcement Learning using evolutionary optimization. We evaluate our genetic variations in the Coin Game environment and also compare them to classical approaches. We showed that our Variational Quantum Circuit approaches perform significantly better compared to a neural network with a similar amount of trainable parameters. Compared to the larger neural network, our approaches archive similar results using $97.88\%$ less parameters.
@@ 1 INTRODUCTION
Reinforcement Learning (RL) itself has made impressive strides, outperforming humans in areas like video games [Badia et al., 2020, Schrittwieser et al., 2019]. Alongside this, quantum technologies are emerging, suggesting faster problem-solving and more efficient training in RL [Harrow and Montanaro, 2017]. However, Quantum Reinforcement Learning (QRL) isn’t without challenges, such as instabilities and vanishing gradients [Franz et al., 2022, Chen et al., 2022]. To address these, researchers have turned to evolutionary optimization methods, as proposed by [Chen et al., 2022], which have shown promising results. With the rising prominence of MARL, combining it with quantum techniques has become a research focal point, leading to the development of Multi-Agent Quantum Reinforcement Learning (MAQRL). In this work, each agent is represented as a Variational Quantum Circuits (VQC). We employ a evolutionary algorithm to optimize the parameters of the circuit. We evaluate different generational evolution strategies and conduct a small scale hyperparameter search for key parameters of the VQC. Our aim is to evaluate MAQRL’s capabilities and compare it to traditional RL methods, using the Coin Game as a benchmark.
In this study, we model each agent using Variational Quantum Circuits (VQC), a promising and adaptable representation in the quantum domain. The inherent flexibility of VQCs allows for the encoding of complex information, making them suitable for representing agent behaviors in diverse environments. To fine-tune these quantum circuits and ensure their optimal performance, we harness the power of an evolutionary algorithm. This algorithm iteratively optimizes the parameters of the VQC, guiding the circuit towards improved decision-making and interactions. While evolutionary algorithms have been traditionally employed in classical domains, their application in the quantum realm offers exciting prospects for efficiently navigating the vast parameter space of VQCs. As part of our experiements, we systematically evaluate multiple generational evolution strategies. By comparing their effectiveness, we aim to identify which strategies most beneficially influence the learning trajectories of the VQCs. Furthermore, recognizing the significance of the VQC’s parameters in determining its behavior and effectiveness, we undertake a small scale hyperparameter search. This search is dedicated to fine-tuning key parameters, ensuring the VQC. Central to our research objectives is the evaluation of Multi-Agent Quantum Reinforcement Learning (MAQRL) and its potential contributions to the field. We are particularly interested in benchmarking MAQRL against established Reinforcement Learning (RL) techniques to learn its advantages and areas of improvement. For a robust and fair assessment, we have chosen the Coin Game, a well-regarded environment in multi-agent research, as our testing ground. In summary our contributions are:
@@
We start in section 2 by explaining the basics of MARL and Evolutionary Optimization. We also give a short introduction to Quantum Computing and VQCs, and mention related studies (section 3). After outlining our methodology (section 4) and experimental setup (section 5), we share the results and implications of our experiments in section 6. We end with a summary and thoughts on next steps for research (section 7).
@@ 2.5 Variational Quantum Circuits
VQC, also known as parameterized quantum circuits, are quantum algorithms that act as function approximators and are trained using a classical optimization process. They are commonly used as a drop-in replacement for Neural Networks for Deep RL [Chen et al., 2022, Schuld et al., 2020, Chen and Goan, 2019, Skolik et al., 2021, Chen, 2022]. A VQC is made up of three stages, as can be seen in Figure 1. First, the classical input is embedded into a quantum state in the State Preperation stage $U(x)$ using superposition. In the Variational Layers stage $V(\theta)$, qubits are then entangled and parameterized for training. Finally, in the Measurement stage, the output of the circuit is measured repeatedly to get the expectation value of each qubit.
@@
Variational Layers The second part of the circuit, referred to as Variational Layers, is made up of repeated single qubit rotations and entanglers (Figure 2, everything within the dashed blue area is repeated $L$ times, where $L$ is the layer count). We use a layer architecture inspired by the circuit-centric classifier design [Schuld et al., 2020] in particular. All of the circuits presented in this paper employ three single qubit rotation gates and CNOT gates as entanglers.
@@ 4 APPROACH
In contrast to previous work, we use VQC rather than neural networks to approximate the value of the agent’s actions. This should mainly demonstrate the improved parameter efficiency, as previously denoted, even applied to complex learning tasks. A VQC consists primarily of three components: the input embedding, the repeated variational layers, and the measurement. Figure 1 depicts the VQC we employ.
@@
To convert the classical data into a quantum state, we use Amplitude Embeddings, represented by $U(x)$ in Figure 2. Caused by the high dimensionality of most state spaces, Amplitude Embeddings are currently the only viable embedding strategy that allow for embedding the whole state information, being able to embed $2^{n_{q}}$ states in $n_{q}$ qubits.
@@ 5.1 Coin Game Environment
As of today, we are in the Noisy Intermediate-Scale Quantum era of quantum computing, where we can simulate only a small number of qubits [Preskill, 2018]. This heavily restricts the amount of data we can embed into the quantum circuit. Therefore, we are limited in our choices for the evaluation environment. We chose the Coin Game environment in its $3\times 3$ gridworld version due to its relatively small observation space. The Coin Game, which was created by [Lerer and Peysakhovich, 2017], is a well-known sequential game for assessing RL strategies. Both the Red and Blue agents in the environment are tasked with collecting coins. Beside the agents there is a single coin placed in the grid that corresponds with one of the agents colors. Figure 3 depicts a exemplary state within the Coin Game.
@@ 6.1 Comparing Generational Evolution Strategies
Examining the average scores depicted in Figure 4, the mutation-only strategy emerges as the best strategy. The crossover strategies showcase similar performance, though the layerwise method frequently achieves marginally superior outcomes. The mutation-only strategy starts with an average reward of $5$, dips slightly below $4$ by the 17th generation, and then steadily rises until the 140th generation. From this point, it fluctuates around a score of $7$. In contrast, the layerwise recombination begins at a lower $3.3$, experiences a rapid ascent until the 30th generation, then stabilizes, eventually reaching an average reward of $6$ by the 123rd generation. This is followed by pronounced fluctuations around this value. The random crossover strategy starts close to the mutation-only at $4.7$, but quickly descends to $3$ by the 17th generation. It then steadily climbs until the 131st generation, achieving a score of $6$. However, this score is not sustained and eventually settles around $5.5$, making it the least effective of the three methods.
Beyond score comparison, we evaluated the average number of coins collected during the experiments. As inferred from the scores, the mutation-only strategy consistently collects more coins, as shown in 5(a). Although there are periods where the strategies yield almost identical coin counts, at other times, a gap of up to $2$ coins is evident. On average, the combined strategies lag slightly behind the mutation-only in terms of coin collection.
The layerwise crossover’s initial surge in Figure 4 correlates with the uptrend in collected coins shown in 5(a) and the coin rate detailed in 5(c). When comparing the strategies based on these metrics, mutation-only consistently achieves the highest coin rate over all generations and collects the most coins, accounting for its superior average reward. The random crossover strategy, while collecting more coins than the layerwise approach, has a significantly reduced coin rate, resulting in diminished overall rewards.
Exploring the coin rate, depicted in 5(c), the layerwise strategy leads until the 90th generation. After that, its rate declines, while the mutation-only strategy exhibits a gradual, consistent rise. A higher coin rate indicates enhanced agent cooperation within the testing environment. This coin rate, combined with the number of coins collected, determines an agent’s reward.
@@ 6.2 Assessing Varying Layer Counts
Inspecting the average rewards in Figure 6, all VQCs, bar the 4-layered one which starts slightly below $3$, commence with scores ranging from $5$ to $5.5$. By the 25th generation, each VQC stabilizes around a reward of $4$. The 4-layer VQC then gradually ascends, consistently holding an average reward of $5$ from the 175th generation. The 6-layer VQC exhibits a steady rise until the 62nd generation, with a more pronounced increase after that, peaking at $6.7$ around the 165th generation and subsequently oscillating around $6.5$. The 8-layer VQC consistently outperforms the others, reaching a reward of $5.5$ by the 70th generation, encountering a brief plateau, and then climbing to $7$ by the 140th generation. The 16-layer VQC, meanwhile, showcases a pronounced growth phase between the 25th and 70th generations, stabilizing around $6$ before another rise to $6.5$ around the 160th generation.
For a comprehensive understanding, we next probe the average coin collection in 7(a). The 4-layer VQC consistently tops the coin collection metric, progressing from just below $7$ to $8$. The 6-layer VQC commences at $6.5$, dips to $5.2$ by the 23rd generation, and then rises to 8 by the 165th generation. The 8-layer VQC, despite securing the highest average reward, begins at $6$ and only stabilizes around $8$ after the 180th generation. The 16-layer VQC, after an initial dip, witnesses a rapid increase from the 24th to 103rd generation, briefly declines, and then fluctuates around $8$ coins. Towards the concluding generations, VQCs with more than 4 layers converge to collect approximately 8 coins.
Analyzing the own coin count in 7(b), we observe that, except for the 4-layer VQC, all VQCs initially decline before ascending. The 4-layer VQC displays a steady yet modest climb, concluding at a count of $6.5$. The 6-layer VQC takes the longest to commence its ascent, eventually oscillating around a count of $7.2$. The 8-layer VQC initiates its climb earlier, achieving a slightly higher count of $7.5$ by the end. The 16-layer VQC, notable for its rÏapid early ascent, consistently hovers around a count of $7$ after the 100th generation. Among the VQCs, the 4-layer variant lags, collecting over one own coin fewer than its counterparts.
@@ 6.3.1 Comparing VQC and Random
First, we compare the results of our VQC approaches to the results of the random baseline. In Figure 8, the score for the random acting agents is approximately 00, since the cooperative sequential coin game is a zero-sum game. The evolutionary-trained VQC approache, however, perform significantly better leading to an average score around $7$. The total coins collected depicted in 9(a) suggests that in contrast to the random agents, the VQC agents successfully learn to collect coins. The own coins collected correlates with number of collected coins (9(b)). In 9(c) we can see that in neither case the cooperation increases over time. In summary, trained agents performed significantly better than random on all metrics, indicating that the training was successful.
@@ 6.3.2 Comparing VQC and Small NN
As depicted in Figure 8, a better result is achieved by the VQC approach compared to random. On this basis, we compare the performance of this VQC approach to that of a neural network with a comparable number of parameters. Here we exploit the higher expressive power of VQCs compared to conventional neural networks [Chen et al., 2022]. Similar to [Chen et al., 2022], we define the expressive power as the capacity to represent particular functions with a constrained number of parameters. Note that the VQC has $148$ parameters (3 * 6 * 8 + 4). The neural network uses two hidden layers with dimension 3 and 4 respectively, resulting in a parameter count of $147$. Both the neural network and the VQC, are trained with mutation only with mutation power $\sigma=0.01$. In Figure 8, we can see that the neural network reward fluctuates in the range of $2.5$ to $3$. As previously discussed in the last section, the VQC approach exhibits a slow learning curve leading to a significant higher score therefore consequently outperforming this neural network. The inferior performance can be explained by the small number of hidden units and parameters present in neural networks. Typically, the number of hidden units is chosen much higher. Further evidence of the neural network’s deficiency is provided by the average number of coins collected. As shown in 9(a), the NN’s number of collected coins is below the average score of the random agents until generation $115$ and after that slightly over it. In comparison, the VQC with the same number of parameters collects two times as many coins on average. The neural network is able to outperform random agents on the basis of its collected own coins, what can be seen in 9(b), leading to the better performance regarding the own coin rate (9(c)). In terms of collected own coins and the own coin rate, the neural network performs significantly worse than the VQC with nearly the same number of parameters. A neural network with this few hidden units and, consequently, parameters is not able learn in the coin game environment successfully. This demonstrates the power of VQCs for RL archiving significantly higher with the same amount of parameters.
@@ 6.3.3 Comparing VQC and Big NN
In the previous section, we observed that a neural network with the same number of parameters as the VQC of our approach cannot match the VQC’s performance. We will now compare the results with a neural network that has significantly more parameters. Again, mutation only is used for the evolution of subsequent generations in both cases and the mutation power is $\sigma=0.01$. We chose a fully connected NN with two hidden layers of size $64$, resulting in a parameter count of $6788$. If we first examine the reward of the neural network and the VQC, we can see in Figure 8 that both produce very similar results over time. Initially, the VQC has a slightly higher score. From generation $50$ onward, there are only minor differences in terms of average score. Overall, the two strategies yield a score of approximatly $7$. On the basis of our experiments, we can say that the VQC achieves nearly identical performance in the Coin Game environment compared to a neural network that has $46$ times more parameters. In 9(a), the initial value of the VQC method is again higher than that of the neural network. Due to a steeper learning curve, the neural network is able to compensate the lower starting value and achieves only a slightly smaller number of collected coins. From there are no discernible differences between the two models. A similar performance of the two approaches can be seen in 9(b), where the average number of own coins collected is shown. In Terms of own coin rate, at first, the VQC archives a slightly higher score. However, at Generation 25, the neural network initially achieves a better own coin rate before being slightly lower between Generation 80 and 162. In the end, the neural network is slightly better in terms of the own coin rate.

### v4

@@ Abstract
Multi-Agent Reinforcement Learning is becoming increasingly more important in times of autonomous driving and other smart industrial applications. Simultaneously a promising new approach to Reinforcement Learning arises using the inherent properties of quantum mechanics, reducing the trainable parameters of a model significantly. However, gradient-based Multi-Agent Quantum Reinforcement Learning methods often have to struggle with barren plateaus, holding them back from matching the performance of classical approaches. While gradient free Quantum Reinforcement Learning methods may alleviate some of these challenges, they too are not immune to the difficulties posed by barren plateaus. We build upon an existing approach for gradient free Quantum Reinforcement Learning and propose three genetic variations with Variational Quantum Circuits for Multi-Agent Reinforcement Learning using evolutionary optimization. We evaluate our genetic variations in the Coin Game environment and also compare them to classical approaches. We showed that our Variational Quantum Circuit approaches perform significantly better compared to a neural network with a similar amount of trainable parameters. Compared to the larger neural network, our approaches archive similar results using $97.88\%$ less parameters.
@@ 1 INTRODUCTION
Reinforcement Learning (RL) itself has made impressive strides, outperforming humans in areas like video games [Badia et al., 2020, Schrittwieser et al., 2019]. Alongside this, quantum technologies are emerging, suggesting faster problem-solving and more efficient training in RL [Harrow and Montanaro, 2017]. However, Quantum Reinforcement Learning (QRL) has it’s challenges, such as instabilities and vanishing gradients [Franz et al., 2022, Chen et al., 2022]. To address these, researchers have turned to evolutionary optimization methods, as proposed by [Chen et al., 2022], which have shown promising results. With the rising prominence of MARL, combining it with quantum techniques has become a research focal point, leading to the development of Multi-Agent Quantum Reinforcement Learning (MAQRL). In this work, each agent is represented as a Variational Quantum Circuits (VQC). We employ a evolutionary algorithm to optimize the parameters of the circuit. We evaluate different generational evolution strategies and conduct a small scale hyperparameter search for key parameters of the VQC. Our aim is to evaluate MAQRL’s capabilities and compare it to traditional RL methods, using the Coin Game as a benchmark.
In this study, we model each agent using Variational Quantum Circuits (VQC), a promising and adaptable representation in the quantum domain. The inherent flexibility of VQCs allows for the encoding of complex information, making them suitable for representing agent behaviors in diverse environments. To fine-tune these quantum circuits and ensure their optimal performance, we harness the power of an evolutionary algorithm. This algorithm iteratively optimizes the parameters of the VQC, guiding the circuit towards improved decision-making and interactions. While evolutionary algorithms have been traditionally employed in classical domains, their application in the quantum realm offers exciting prospects for efficiently navigating the vast parameter space of VQCs. As part of our experiements, we systematically evaluate multiple generational evolution strategies. By comparing their effectiveness, we aim to identify which strategies most beneficially influence the learning trajectories of the VQCs. Furthermore, recognizing the significance of the VQC’s parameters in determining its behavior and effectiveness, we undertake a small scale hyperparameter search. This search is dedicated to fine-tuning key parameters, ensuring the VQC. Central to our research objectives is the evaluation of MAQRL and its potential contributions to the field. We are particularly interested in benchmarking MAQRL against established RL techniques to learn its advantages and areas of improvement. For a robust and fair assessment, we have chosen the Coin Game, a well-regarded environment in multi-agent research, as our testing ground. In summary our contributions are:
@@
We start in Section 2 by explaining the basics of MARL and Evolutionary Optimization. We also give a short introduction to Quantum Computing and VQCs, and mention related studies (Section 3). After outlining our methodology (Section 4) and experimental setup (Section 5), we share the results and implications of our experiments in Section 6. We end with a summary and thoughts on next steps for research (Section 7). All code and experiments can be found here [^1] https://github.com/michaelkoelle/qmarl-evo.
@@ 2.5 Variational Quantum Circuits
VQC, also known as parameterized quantum circuits, are quantum algorithms that act as function approximators and are trained using a classical optimization process. They are commonly used as a drop-in replacement for Neural Networks for Deep RL [Chen et al., 2022, Schuld et al., 2020, Chen and Goan, 2019, Skolik et al., 2021, Chen, 2022]. A VQC is made up of three stages, as can be seen in Fig. 1. First, the classical input is embedded into a quantum state in the State Preperation stage $U(x)$ using superposition. In the Variational Layers stage $V(\theta)$, qubits are then entangled and parameterized for training. Finally, in the Measurement stage, the output of the circuit is measured repeatedly to get the expectation value of each qubit.
@@
Variational Layers The second part of the circuit, referred to as Variational Layers, is made up of repeated single qubit rotations and entanglers (Fig. 2, everything within the dashed blue area is repeated $L$ times, where $L$ is the layer count). We use a layer architecture inspired by the circuit-centric classifier design [Schuld et al., 2020] in particular. All of the circuits presented in this paper employ three single qubit rotation gates and CNOT gates as entanglers.
@@ 4 APPROACH
In contrast to previous work, we use VQC rather than neural networks to approximate the value of the agent’s actions. This should mainly demonstrate the improved parameter efficiency, as previously denoted, even applied to complex learning tasks. A VQC consists primarily of three components: the input embedding, the repeated variational layers, and the measurement. Fig. 1 depicts the VQC we employ.
@@
To convert the classical data into a quantum state, we use Amplitude Embeddings, represented by $U(x)$ in Fig. 2. Caused by the high dimensionality of most state spaces, Amplitude Embeddings are currently the only viable embedding strategy that allow for embedding the whole state information, being able to embed $2^{n_{q}}$ states in $n_{q}$ qubits.
@@ 5.1 Coin Game Environment
As of today, we are in the Noisy Intermediate-Scale Quantum era of quantum computing, where we can simulate only a small number of qubits [Preskill, 2018]. This heavily restricts the amount of data we can embed into the quantum circuit. Therefore, we are limited in our choices for the evaluation environment. We chose the Coin Game environment in its $3\times 3$ gridworld version due to its relatively small observation space. The Coin Game, which was created by [Lerer and Peysakhovich, 2017], is a well-known sequential game for assessing RL strategies. Both the Red and Blue agents in the environment are tasked with collecting coins. Beside the agents there is a single coin placed in the grid that corresponds with one of the agents colors. Fig. 3 depicts a exemplary state within the Coin Game.
@@ 6.1 Comparing Generational Evolution Strategies
Examining the average scores depicted in Fig. 4, the mutation-only strategy emerges as the best strategy. The crossover strategies showcase similar performance, though the layerwise method frequently achieves marginally superior outcomes. The mutation-only strategy starts with an average reward of $5$, dips slightly below $4$ by the 17th generation, and then steadily rises until the 140th generation. From this point, it fluctuates around a score of $7$. In contrast, the layerwise recombination begins at a lower $3.3$, experiences a rapid ascent until the 30th generation, then stabilizes, eventually reaching an average reward of $6$ by the 123rd generation. This is followed by pronounced fluctuations around this value. The random crossover strategy starts close to the mutation-only at $4.7$, but quickly descends to $3$ by the 17th generation. It then steadily climbs until the 131st generation, achieving a score of $6$. However, this score is not sustained and eventually settles around $5.5$, making it the least effective of the three methods.
Beyond score comparison, we evaluated the average number of coins collected during the experiments. As inferred from the scores, the mutation-only strategy consistently collects more coins, as shown in Fig. 5(a). Although there are periods where the strategies yield almost identical coin counts, at other times, a gap of up to $2$ coins is evident. On average, the combined strategies lag slightly behind the mutation-only in terms of coin collection.
The layerwise crossover’s initial surge in Fig. 4 correlates with the uptrend in collected coins shown in Fig. 5(a) and the coin rate detailed in Fig. 5(c). When comparing the strategies based on these metrics, mutation-only consistently achieves the highest coin rate over all generations and collects the most coins, accounting for its superior average reward. The random crossover strategy, while collecting more coins than the layerwise approach, has a significantly reduced coin rate, resulting in diminished overall rewards.
Exploring the coin rate, depicted in Fig. 5(c), the layerwise strategy leads until the 90th generation. After that, its rate declines, while the mutation-only strategy exhibits a gradual, consistent rise. A higher coin rate indicates enhanced agent cooperation within the testing environment. This coin rate, combined with the number of coins collected, determines an agent’s reward.
@@ 6.2 Assessing Varying Layer Counts
Inspecting the average rewards in Fig. 6, all VQCs, bar the 4-layered one which starts slightly below $3$, commence with scores ranging from $5$ to $5.5$. By the 25th generation, each VQC stabilizes around a reward of $4$. The 4-layer VQC then gradually ascends, consistently holding an average reward of $5$ from the 175th generation. The 6-layer VQC exhibits a steady rise until the 62nd generation, with a more pronounced increase after that, peaking at $6.7$ around the 165th generation and subsequently oscillating around $6.5$. The 8-layer VQC consistently outperforms the others, reaching a reward of $5.5$ by the 70th generation, encountering a brief plateau, and then climbing to $7$ by the 140th generation. The 16-layer VQC, meanwhile, showcases a pronounced growth phase between the 25th and 70th generations, stabilizing around $6$ before another rise to $6.5$ around the 160th generation.
For a comprehensive understanding, we next probe the average coin collection in Fig. 7(a). The 4-layer VQC consistently tops the coin collection metric, progressing from just below $7$ to $8$. The 6-layer VQC commences at $6.5$, dips to $5.2$ by the 23rd generation, and then rises to 8 by the 165th generation. The 8-layer VQC, despite securing the highest average reward, begins at $6$ and only stabilizes around $8$ after the 180th generation. The 16-layer VQC, after an initial dip, witnesses a rapid increase from the 24th to 103rd generation, briefly declines, and then fluctuates around $8$ coins. Towards the concluding generations, VQCs with more than 4 layers converge to collect approximately 8 coins.
Analyzing the own coin count in Fig. 7(b), we observe that, except for the 4-layer VQC, all VQCs initially decline before ascending. The 4-layer VQC displays a steady yet modest climb, concluding at a count of $6.5$. The 6-layer VQC takes the longest to commence its ascent, eventually oscillating around a count of $7.2$. The 8-layer VQC initiates its climb earlier, achieving a slightly higher count of $7.5$ by the end. The 16-layer VQC, notable for its rÏapid early ascent, consistently hovers around a count of $7$ after the 100th generation. Among the VQCs, the 4-layer variant lags, collecting over one own coin fewer than its counterparts.
@@ 6.3.1 Comparing VQC and Random
First, we compare the results of our VQC approaches to the results of the random baseline. In Fig. 8, the score for the random acting agents is approximately 00, since the cooperative sequential coin game is a zero-sum game. The evolutionary-trained VQC approache, however, perform significantly better leading to an average score around $7$. The total coins collected depicted in Fig. 9(a) suggests that in contrast to the random agents, the VQC agents successfully learn to collect coins. The own coins collected correlates with number of collected coins (Fig. 9(b)). In Fig. 9(c) we can see that in neither case the cooperation increases over time. In summary, trained agents performed significantly better than random on all metrics, indicating that the training was successful.
@@ 6.3.2 Comparing VQC and Small NN
As depicted in Fig. 8, a better result is achieved by the VQC approach compared to random. On this basis, we compare the performance of this VQC approach to that of a neural network with a comparable number of parameters. Here we exploit the higher expressive power of VQCs compared to conventional neural networks [Chen et al., 2022]. Similar to [Chen et al., 2022], we define the expressive power as the capacity to represent particular functions with a constrained number of parameters. Note that the VQC has $148$ parameters (3 * 6 * 8 + 4). The neural network uses two hidden layers with dimension 3 and 4 respectively, resulting in a parameter count of $147$. Both the neural network and the VQC, are trained with mutation only with mutation power $\sigma=0.01$. In Fig. 8, we can see that the neural network reward fluctuates in the range of $2.5$ to $3$. As previously discussed in the last section, the VQC approach exhibits a slow learning curve leading to a significant higher score therefore consequently outperforming this neural network. The inferior performance can be explained by the small number of hidden units and parameters present in neural networks. Typically, the number of hidden units is chosen much higher. Further evidence of the neural network’s deficiency is provided by the average number of coins collected. As shown in Fig. 9(a), the NN’s number of collected coins is below the average score of the random agents until generation $115$ and after that slightly over it. In comparison, the VQC with the same number of parameters collects two times as many coins on average. The neural network is able to outperform random agents on the basis of its collected own coins, what can be seen in Fig. 9(b), leading to the better performance regarding the own coin rate (Fig. 9(c)). In terms of collected own coins and the own coin rate, the neural network performs significantly worse than the VQC with nearly the same number of parameters. A neural network with this few hidden units and, consequently, parameters is not able learn in the coin game environment successfully. This demonstrates the power of VQCs for RL archiving significantly higher with the same amount of parameters.
@@ 6.3.3 Comparing VQC and Big NN
In the previous section, we observed that a neural network with the same number of parameters as the VQC of our approach cannot match the VQC’s performance. We will now compare the results with a neural network that has significantly more parameters. Again, mutation only is used for the evolution of subsequent generations in both cases and the mutation power is $\sigma=0.01$. We chose a fully connected NN with two hidden layers of size $64$, resulting in a parameter count of $6788$. If we first examine the reward of the neural network and the VQC, we can see in Fig. 8 that both produce very similar results over time. Initially, the VQC has a slightly higher score. From generation $50$ onward, there are only minor differences in terms of average score. Overall, the two strategies yield a score of approximatly $7$. On the basis of our experiments, we can say that the VQC achieves nearly identical performance in the Coin Game environment compared to a neural network that has $46$ times more parameters. In Fig. 9(a), the initial value of the VQC method is again higher than that of the neural network. Due to a steeper learning curve, the neural network is able to compensate the lower starting value and achieves only a slightly smaller number of collected coins. From there are no discernible differences between the two models. A similar performance of the two approaches can be seen in Fig. 9(b), where the average number of own coins collected is shown. In Terms of own coin rate, at first, the VQC archives a slightly higher score. However, at Generation 25, the neural network initially achieves a better own coin rate before being slightly lower between Generation 80 and 162. In the end, the neural network is slightly better in terms of the own coin rate.